OpenStudy (anonymous):

The expression (secx + tanx)2 is the same as _____. A. 1 + 2cscx B. sec2x + 2cscx + tan2x C. sec2x + tan2x D. 1 + 2tan2x + 2secx tanx

zepdrix (zepdrix):

\[\Large\rm (\sec x + \tan x)^2\]Remember how to expand out a squared expression like this? :)\[\Large\rm (a+b)^2=a^2+2ab+b^2\]

zepdrix (zepdrix):

Hmm so judging by the options, you'll have to simplify a lil bit after expanding...

OpenStudy (anonymous):

thank you

zepdrix (zepdrix):

If you're able to get past that first part, here is a hint to help you finish it up. Remember your Pythagorean identity for tangent,\[\Large\rm \color{red}{\sec^2x=1+\tan^2x}\]

OpenStudy (anonymous):

pythagorean identity for tan?

zepdrix (zepdrix):

These are the Pythagorean Identities:\[\Large\rm \sin^2x+\cos^2x=1\]\[\Large\rm 1+\tan^2x=\sec^2x\]\[\Large\rm 1+\cot^2x=\csc^2x\]

zepdrix (zepdrix):

After you expand everything out, you can replace the sec^2x(that you'll end up with) with 1+tan^2x.

OpenStudy (anonymous):

so i would get 1+tan^2x=1+tan^2x? then what?

OpenStudy (anonymous):

i think i did something wrong there

OpenStudy (anonymous):

could you please walk me through the whole thing?

OpenStudy (anonymous):

oh wait i get it!

OpenStudy (anonymous):

sorry just got off a flight and im trying to work while i wait for the next one

OpenStudy (anonymous):

anyways thank you very much!

zepdrix (zepdrix):

Think you got it? c: cool

OpenStudy (anonymous):

definitely :) there is another question i was wondering if you could help me with though

zepdrix (zepdrix):

Sure

OpenStudy (anonymous):

Which of the following is not an identity? A. secx cscx(tanx + cotx) = sec2x + csc2x B. sin2x + tan2x + cos2x = sec2x C. 2sin2x - sinx = 1 D. 2cos2x - 1 = 1 - 2sin2x

OpenStudy (anonymous):

I'm not so great with identities

zepdrix (zepdrix):

Are any of those angles (2x) or are they all squares? It's kind of hard to tell the difference when you copy paste :P heh

zepdrix (zepdrix):

For example is option C:\[\Large\rm 2\sin^2x-\sin x=1, \qquad\text{or}\qquad 2\sin(2x)-\sin x=1\]

OpenStudy (anonymous):

they're all squares, sorry

zepdrix (zepdrix):

Let's look at option B a moment: \[\Large\rm \sin^2x + \tan^2x + \cos^2x = \sec^2x\]Let's move things around a lil bit,\[\Large\rm \color{royalblue}{\sin^2x + \cos^2x} + \tan^2x = \sec^2x\]

zepdrix (zepdrix):

Hmm do you see an identity we can apply to the blue part? Gotta remember those Pythagorean Identities!! :) Scroll back up the page if you forgot them.

OpenStudy (anonymous):

\[\sin^2x+\cos^2x=1\]

OpenStudy (anonymous):

and then 1+tan^2x=sec^2x

zepdrix (zepdrix):

Ok great! So option B is not what we're looking for.

OpenStudy (anonymous):

alrighty, on to the next one

zepdrix (zepdrix):

Option D:\[\Large\rm 2\cos^2x - 1 \quad=\quad 1 - 2\sin^2x\]Hmm let's see...

zepdrix (zepdrix):

Let's use our Pythagorean Identity to substitute something in for cos^2x.\[\Large\rm \sin^2x+\cos^2x=1\qquad\to\qquad \color{orangered}{\cos^2x=1-\sin^2x}\] We'll use that to replace our cos^2x,\[\Large\rm 2(\color{orangered}{\cos^2x}) - 1 \quad=\quad 1 - 2\sin^2x\]

zepdrix (zepdrix):

Understand what I did with the Pythagorean Identity? I solved it for cos^2x.

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

Plugging in gives us, \[\Large\rm 2(\color{orangered}{1-\sin^2x}) - 1 \quad=\quad 1 - 2\sin^2x\] Hmm will it simplify to match the other side or no?

OpenStudy (anonymous):

yes it will

OpenStudy (anonymous):

i think

zepdrix (zepdrix):

Ok cool c: So we don't want option D!

OpenStudy (anonymous):

so it will be c or a

zepdrix (zepdrix):

This is a tough problem, it's really tedious, and requires you to know a lot of your identities.

zepdrix (zepdrix):

Umm ok let's check out option A:\[\Large\rm \sec x \csc x (\tan x + \cot x) = \sec^2x + \csc^2x\]

zepdrix (zepdrix):

In order to work with this one, I think we're going to want to convert everything to sines and cosines.

OpenStudy (anonymous):

alright

zepdrix (zepdrix):

Remember how to write secx, cscx, tanx and cotx in terms of sines and cosines?

OpenStudy (anonymous):

no :(

zepdrix (zepdrix):

You... <.<

OpenStudy (anonymous):

Hahaha sorry, this is my math prep for precalc next year, i barely know anything

zepdrix (zepdrix):

\[\Large\rm \sec x =\frac{1}{\cos x},\qquad \csc x=\frac{1}{\sin x}\] \[\Large\rm \tan x=\frac{\sin x}{\cos x},\qquad \cot x=\frac{\cos x}{\sin x}\]

zepdrix (zepdrix):

\[\Large\rm \sec x \csc x (\tan x + \cot x) = \sec^2x + \csc^2x\] Plugging all those identities in gives us:\[\Large\rm \frac{1}{\sin x \cos x} \left(\frac{\sin x}{\cos x} + \frac{\cos x}{\sin x}\right) = \sec^2x + \csc^2x\]Ok? :o

OpenStudy (anonymous):

Okay, that makes sense

zepdrix (zepdrix):

We'll distribute the ugly fraction in front, to each term in the brackets.

zepdrix (zepdrix):

\[\Large\rm \frac{\sin x}{\sin x \cos x\cos x} + \frac{\cos x}{\sin x\sin x\cos x} = \sec^2x + \csc^2x\]

zepdrix (zepdrix):

Understand what I did there? :o I gave them each a 1/(sin x cos x)

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

Looks like we have some nice cancellations. \[\Large\rm \frac{\cancel{\sin x}}{\cancel{\sin x} \cos x\cos x} + \frac{\cancel{\cos x}}{\sin x\sin x\cancel{\cos x}} = \sec^2x + \csc^2x\]

zepdrix (zepdrix):

\[\Large\rm \frac{1}{\cos^2x}+\frac{1}{\sin^2x}=\sec^2x+\csc^2x\]

zepdrix (zepdrix):

Is there a way we can write these back in terms of secant and cosecant? Look back at the 4 identities I posted before! :O

OpenStudy (anonymous):

it would be secx and cscx?

zepdrix (zepdrix):

Lemme write the squares like this so maybe it's easier to see what's going on,\[\Large\rm \left(\frac{1}{\cos x}\right)^2+\left(\frac{1}{\sin x}\right)^2=\sec^2x+\csc^2x\]

zepdrix (zepdrix):

We had an identity for 1/cos x, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\rm \left(\color{royalblue}{\frac{1}{\cos x}}\right)^2+\left(\color{royalblue}{\frac{1}{\sin x}}\right)^2=\sec^2x+\csc^2x\]Plugging those identities in gives us,\[\Large\rm \left(\color{royalblue}{\sec x}\right)^2+\left(\color{royalblue}{\csc x}\right)^2=\sec^2x+\csc^2x\]

zepdrix (zepdrix):

So is this one true? :O

OpenStudy (anonymous):

i think so yes

zepdrix (zepdrix):

Ok good! So it looks like we've narrowed it down to C! I didn't mean to check every correct one first, that was on accident lol

OpenStudy (anonymous):

hahaha it's fine :) it helped me understand the process! thank you so much!

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

woohoo :)